The regular price for a pair of shoes is $245. They are currently on sale for 25% off.  The harmonized sales tax in Ontario is 13%. Your receipt shows that the final cost is $ 215.32. Are you happy with the final cost? Explain why or why not.  Complete without a calculator and show your work.


1
Expert's answer
2022-05-23T12:15:50-0400

The regular price for a pair of shoes is $245. They are currently on sale for 25% off.


"\$245(1-0.25)=\$245\cdot\dfrac{3}{4}"

The harmonized sales tax in Ontario is 13%


"\$245\cdot\dfrac{3}{4}\cdot(1+0.13)<\$245\cdot\dfrac{3.4}{4}"

"=\$245\cdot0.85=\$208.25"

We have to pay less than "\$208.25." But my receipt shows that the final cost is $ 215.32


Since "\$208.25<\$215.32" I don't agree with the final cost $ 215.32 . I am not happy with the final cost.